OpenStudy (anonymous):

Consuelo is shopping and sees that the price of 4 melons is $7.00. write a proportion that Consuelo can use to find the price of 1 melon

OpenStudy (anonymous):

Hi @xokatexo - so, do you know how to start this problem? How to put it in an equation?

OpenStudy (anonymous):

no thats what i need help finding. I can solve it from theere

OpenStudy (anonymous):

No problem! So, let's let x stand for melons. Therefore, we have: \[4x=7.00\]

OpenStudy (anonymous):

i actually need a proportion. sorry

OpenStudy (anonymous):

So, you need a proportion, correct? Now, you know that 4 melons equals $7.00, and you want a proportion that will allow you to find one melon

OpenStudy (anonymous):

wait i think i got it

OpenStudy (anonymous):

hehe, beat me to it ;) So, a proportion would look like this: \[4x/7.00 = x/y\]

OpenStudy (anonymous):

is it 4/7=1/x

OpenStudy (anonymous):

where y equals the cost

OpenStudy (anonymous):

Yup, you got it! (I think) Let me check with someone real quick

OpenStudy (anonymous):

who are you checking with?

OpenStudy (anonymous):

@Diyadiya - got any input? Is \[4/7=1/x\] correct?

OpenStudy (anonymous):

yeah i got 1.75 as my final answer

OpenStudy (diyadiya):

4melons = 7$ 1melon= 7/4$ ?

OpenStudy (diyadiya):

Yes that's right :D you found out 1/x=4/7 so x =7/4 Yes thats right :)